# Spectral Convolutional Conditional Neural Processes

### v1

@@ 1 Introduction
Since their initial introduction, there have been various follow-up works aimed at addressing the limitations of vanilla CNPs, including issues such as underfitting (Kim et al., 2019; Nguyen & Grover, 2022; Kim et al., 2022), uncorrelated predictions (Louizos et al., 2019; Foong et al., 2020; Bruinsma et al., 2021; Markou et al., 2022; Bruinsma et al., 2023), and the use of a simple Gaussian predictive distribution (mohseni23a). In this article, we focus specifically on Convolutional Conditional Neural Processes (ConvCNPs, Gordon et al. (2019)), which represent a variant designed to handle symmetries such as translation equivariance by integrating convolutional neural networks (CNNs). Although ConvCNPs exhibit impressive performance, they may encounter challenges when it comes to effectively integrating partial information scattered irregularly across the domain to infer the underlying pattern. This arises from the commonly used local discrete kernels, which often possess an effective memory horizon considerably shorter than the length of the input sequence (Romero et al., 2021). To circumvent this issue, we take a different approach based on the developments of neural network based surrogates for solving partial differential equations (PDEs, Raissi et al. (2019); Lu et al. (2019); Kovachki et al. (2023)). In particular, we leverage the formulation of Fourier Neural Operators (FNOs, Li et al. (2020a)) to perform global convolution. The resulting model, which we refer to as Spectral Convolution Conditional Neural Process (SConvCNP), is demonstrated to outperform baselines on challenging regression problems [^1] Code is available at https://github.com/peiman-m/SpecConvCNP.
@@ 3.1 Convolutional Conditional Neural Processes
ConvCNP
SConvCNP

### v2

@@ 1 Introduction
Since their initial introduction, there have been various follow-up works aimed at addressing the limitations of vanilla CNPs, including issues such as underfitting (Kim et al., 2019; Nguyen & Grover, 2022; Kim et al., 2022), uncorrelated predictions (Louizos et al., 2019; Foong et al., 2020; Bruinsma et al., 2021; Markou et al., 2022; Bruinsma et al., 2023), and the use of a simple Gaussian predictive distribution (Mohseni et al., 2023). In this article, we focus specifically on Convolutional Conditional Neural Processes (ConvCNPs, Gordon et al. (2019)), which represent a variant designed to handle symmetries such as translation equivariance by integrating convolutional neural networks (CNNs). Although ConvCNPs exhibit impressive performance, they may encounter challenges when it comes to effectively integrating partial information scattered irregularly across the domain to infer the underlying pattern. This arises from the commonly used local discrete kernels, which often possess an effective memory horizon considerably shorter than the length of the input sequence (Romero et al., 2021). To circumvent this issue, we take a different approach based on the developments of neural network based surrogates for solving partial differential equations (PDEs, Raissi et al. (2019); Lu et al. (2019); Kovachki et al. (2023)). In particular, we leverage the formulation of Fourier Neural Operators (FNOs, Li et al. (2020a)) to perform global convolution. The resulting model, which we refer to as Spectral Convolution Conditional Neural Process (SConvCNP), is demonstrated to outperform baselines on challenging regression problems [^1] Code is available at https://github.com/peiman-m/SpecConvCNP.
@@ 3.1 Convolutional Conditional Neural Processes
ConvCNP Refer to caption Refer to caption Refer to caption
SConvCNP Refer to caption Refer to caption Refer to caption